### v1

# Non-linear shattered pellet injection modelling in ASDEX Upgrade

Abstract
Shattered pellet injection (SPI) is selected for the disruption mitigation system in ITER, due to deeper penetration, expected assimilation efficiency and prompt material delivery. This article describes non-linear simulations of SPI in the ASDEX Upgrade tokamak to test the mitigation efficiency of different injection parameters for neon-doped deuterium pellets using the JOREK code. The simulations are executed as fluid simulations. Additional marker particles are used to evolve the charge state distribution of impurities based on OpenADAS atomic data, i.e., no coronal equilibrium assumption is made. Regarding the pellet composition, neon fraction scans from 0 - 10% are performed. Numerical results show that the thermal quench (TQ) occurs in two stages. In the first stage, approximately half of the thermal energy is lost abruptly, primarily through convective and conductive transport in the stochastic fields. This stage is relatively independent of the injection parameters. In the second stage, where the majority of the remaining thermal energy is lost, radiation plays a dominant role. In cases of very low neon content, this second stage may not occur at all. A larger fraction ($\sim$20%) of the total material in the pellet is assimilated in the plasma for low neon fraction pellets (0.12%) since the full thermal collapse of the plasma occurs later than in high neon fraction scenarios. Nevertheless, the total number of assimilated neon atoms increases with increasing neon fraction. The effects of fragment size and penetration speed are further studied. Slower and smaller fragments promote edge cooling and the formation of a cold front. Faster fragments result in shorter TQ duration and higher assimilation as they reach the hotter plasma regions quicker. Using synthetic diagnostics, comparisons of general trend between simulations and experiments are conducted.
Keywords: tokamak, MHD, disruption, shattered pellet injection
[…]
1 Introduction
Unmitigated major disruptions [1] could potentially damage the plasma facing components and the structural integrity of large fusion devices due to heat loads, electromagnetic forces, and runaway electrons (REs) [2]. To effectively mitigate all these above consequences, the shattered pellet injection (SPI) is selected as the disruption mitigation system (DMS) of ITER [3]. SPI is a technology that can actively shut down the ‘unhealthy’ plasma, by injecting a cryogenic pellet composed of pure deuterium or a mixture containing high-Z impurities [4, 5]. The injected deuterium aids in suppressing REs by increasing collisions, thereby decelerating these fast electrons, while the ablated impurities can radiate away the stored energy uniformly, hence reducing the damage to the plasma facing components. To develop a thorough understanding of SPI, essential for extrapolating its performance to future machines like ITER, SPI systems are installed on tokamak devices worldwide, including DIII-D [6], JET [7], KSTAR [8], J-TEXT [9], HL-2A [10], EAST [11], and ASDEX Upgrade (AUG) [12]. On the modelling side, many efforts are being undertaken to study the material assimilation, RE mitigation and magnetohydrodynamic (MHD) response after injection. Various codes, including the 3D MHD codes NIMROD [13], M3D-C1 [14] and JOREK [15], as well as lower dimensional codes like INDEX [16, 17] and DREAM [18], are being employed in these studies.
To validate the optimum fragment size, speed, and neon mixture fraction for the ITER DMS, the AUG tokamak has installed a flexible SPI system with three independent injecting lines, allowing for multiple simultaneous injections with options of different shatter heads [19]. After extensive lab tests and commissioning [12, 19], SPI experiments with different shattering scenarios were carried out in AUG during the 2021/2022 campaign [20]. Due to precise lab characterization of the shard cloud [21] and an excellent bolometry coverage [20], experiments in AUG and simulations for AUG scenarios can advance the understanding of the involved processes substantially and can help to robustly validate the models in view of reliable predictive simulations for ITER.
The 3D non-linear studies addressed here aim to gain a deeper understanding of complex physical phenomena by fully accounting for various self-consistent processes that are simplified or neglected in lower-dimensional models. These studies, although computationally more expensive, are crucial because they allow for the detailed inclusion of more physical processes, such as relaxation of the temperature profile by stochastic transport, current profile evolution due to resistive decay in the presence of the very large temperature variations [22], and plasmoid drift dynamics originating in the vertical motion of the electrons and ions in the inhomogeneous magnetic field [23, 24].
The detailed nature of 3D non-linear studies also allows for direct comparisons with experimental data via synthetic diagnostics, e.g., regarding localized density measurements [25, 26], wall heat deposition [27], and electromagnetic forces in a 3D context [28], which are essential for understanding real-world scenarios. However, despite the advancements these studies provide, there is still much work to be done in interpreting experimental results and validating these models against empirical data. This is where systematic parameter scans become invaluable, especially given the challenges associated with full thermal quench (TQ) simulations.
In the present article, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using the JOREK code [29, 30]. Due to limitation of computational resource and numerical difficulty, our simulations are aimed for full TQ to early current quench (CQ) phase. Mitigation efficiency of different injection parameters for neon-doped deuterium pellets is numerically studied. Detailed analysis and variations of the injection parameters reveal fundamental dependencies. The rest of the paper will be structured as follows: In section 2, the governing reduced MHD equations are introduced, as well as the shards configuration describing the shattering process. In section 3, the effectiveness of SPI under various injection parameters are investigated. Finally, the paper is summarized and conclusions are drawn in section 4.
2 Simulation model
2.1 The governing equations
For our studies we employ the visco-resistive two-temperature reduced MHD model of JOREK in the realistic tokamak geometry as briefly summarized in the following.
$\frac{\partial\psi}{\partial t}=\eta\Delta^{*}\psi-R\{u,\psi\}-F_{0}\frac{\partial u}{\partial\phi},$ (1)
$R\nabla$ $\cdot\left[R^{2}\left(\rho\nabla_{pol}\frac{\partial u}{\partial t}+\nabla_{pol}u\frac{\partial\rho}{\partial t}\right)\right]=$ (2)
$\frac{1}{2}\left\{R^{2}\left|\nabla_{pol}u\right|^{2},R^{2}\rho\right\}+\left\{R^{4}\rho\omega,u\right\}$
$-R\nabla\cdot\left[R^{2}\nabla_{pol}u\nabla\cdot(\rho\mathbf{v})\right]+\{\psi,j\}-\frac{F_{0}}{R}\frac{\partial j}{\partial\phi}$
+$+\left\{P,R^{2}\right\}+R\mu\nabla_{pol}^{2}\omega,$
$B^{2}\frac{\partial}{\partial t}\left(\rho v_{\|}\right)=$ $-\frac{1}{2}\rho\frac{F_{0}}{R^{2}}\frac{\partial}{\partial\phi}\left(v_{\|}B\right)^{2}-\frac{\rho}{2R}\left\{B^{2}v_{\|}^{2},\psi\right\}$ (3)
$-\frac{F_{0}}{R^{2}}\frac{\partial P}{\partial\phi}+\frac{1}{R}\{\psi,P\}$
$-B^{2}\nabla\cdot(\rho\mathbf{v})v_{\|}+B^{2}\mu_{\|}\nabla_{pol}^{2}v_{\|},$
$\frac{\partial}{\partial t}\rho=$ $-\nabla\cdot(\rho\mathbf{v})+\nabla\cdot\left[D_{D}\nabla\left(\rho-\rho_{imp}\right)\right]$ (4)
$+\nabla\cdot\left(D_{imp}\nabla\rho_{imp}\right)+S_{D}+S_{imp},$
$\frac{\partial}{\partial t}\rho_{imp}=-\nabla\cdot\left(\rho_{imp}\mathbf{v}\right)+\nabla\cdot\left(D_{imp}\nabla\rho_{imp}\right)+S_{imp},$ (5)
$\frac{\partial}{\partial t}P_{i}=$ $-\mathbf{v}\cdot\nabla P_{i}-\gamma P_{i}\nabla\cdot\mathbf{v}$ (6)
$+\nabla\cdot\left(\kappa_{\perp}\nabla_{\perp}T_{i}+\kappa_{i,\|}\nabla_{\|}T_{i}\right)$
$+\frac{\gamma-1}{2}\mathbf{v}\cdot\mathbf{v}\left(S_{D}+S_{imp}\right)+(\gamma-1)\mu_{\|}$
×$\times\left[\nabla_{pol}\left(v_{\|}B\right)\right]^{2}+\left(n_{D}+n_{imp}\right)\left(\partial_{t}T_{i}\right)_{c,e},$
$\frac{\partial}{\partial t}P_{e}=$ $-\mathbf{v}\cdot\nabla P_{e}-\gamma P_{e}\nabla\cdot\mathbf{v}$ (7)
$+\nabla\cdot\left(\kappa_{\perp}\nabla_{\perp}T_{e}+\kappa_{e,\|}\nabla_{\|}T_{e}\right)$
$+\frac{\gamma-1}{R^{2}}\eta j^{2}-(\gamma-1)\left(P_{rad}+P_{ion}\right)$
$+n_{e}\left(\partial_{t}T_{e}\right)_{c,i}.$
The main scalar fields of interest in the reduced MHD system are, the poloidal magnetic flux $\psi$, the current density $j$, the vorticity $\omega$, the stream function $u$, the parallel velocity $v_{\|}$, the total mass density $\rho$, the impurity mass density $\rho_{imp}$, the ion/electron pressure $P_{i/e}$ and the ion/electron temperature $T_{i/e}$. The diffusive coefficients $\eta$, $\mu$, $\mu_{\|}$, $D_{D}$, $D_{imp}$, $\kappa_{\perp}$, $\kappa_{i,\|}$, $\kappa_{e,\|}$ are respectively the resistivity, perpendicular viscosity, parallel viscosity, particle diffusivity for main ions, particle diffusivity for impurity, perpendicular thermal conductivity, parallel thermal conductivity for ions and electrons. The source terms $S_{D}$ and $S_{imp}$ are the SPI ablation sources for main ion and impurity species, respectively. The Poisson bracket is defined as $\{f,g\}=R(\nabla f\times\nabla g)\cdot\nabla\phi$. Detailed simulation parameters are given in section 3.1.
A coronal equilibrium distribution of impurity charge states had been utilized in previous JOREK SPI simulations [31]. To obtain a more accurate evolution of charge states and radiation, the impurity charge states are evolved in time self-consistently using OpenADAS atomic coefficients [32]. Here, marker particles are employed to carry the charge state evolution, while all other quantities are treated in the fluid picture. Markers are created according to the impurity ablation source $S_{imp}$ and are convected with the fluid velocity field. The ionization and recombination processes are calculated according to the local temperature and density fields. The atomic processes taking into account the charge state distribution carried by the makers provides the necessary information to calculate the radiation power density $P_{rad}$ and ionization power density $P_{ion}$.
Table 1: The SPI injection parameters for the different simulations performed in this study. Naming convention: Large/Medium/Small Fragment as well as Full/Half Velocity and the neon atomic mixture ratio in the pellet. * for the half speed case (222 m/s), to isolate the effect of the penetration speed, we maintain the same shard sizes as in LF_FV_Ne10 and MF_FV_Ne10, while manually reducing the velocities of the shards to half of their original values. It is equivalent to keeping the $v_{\perp}$ while reducing the $v_{\parallel}$ of the pellet, such that the LF_HV_Ne10 is comparable to a case with 222 m/s and 25 deg. shatter angle.
Name Ne atoms D atoms Ne frac. Speed Shatter angle Frag. Num.
LF_FV_Ne0 0 $3.0\times 10^{22}$ 0% 443 m/s 12.5 deg. 53
LF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 12.5 deg. 53
LF_FV_Ne1 $3.0\times 10^{20}$ $3.0\times 10^{22}$ 1% 443 m/s 12.5 deg. 53
LF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 12.5 deg. 53
MF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 20.0 deg. 199
SF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 25.0 deg. 1105
MF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 20.0 deg. 199
SF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 25.0 deg. 1105
LF_HV_Ne10* $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 222 m/s 25.0 deg. 53
MF_HV_Ne10* $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 222 m/s N/A 199
Figure 1: Shards plumes generated from the 443 m/s pellets. From left to right, the shattering angles are set as $12.5^{\circ}$ (GT3 in AUG), $20^{\circ}$ and $25^{\circ}$ (GT1 in AUG), respectively. The produced shard numbers are accordingly 53, 199 and 1105. The contours indicate the locations of different equilibrium resonant surfaces.
Figure 2: Equilibrium profiles of safety factor $q$, electron density $n_{e}$, toroidal current density $j$ and electron temperature $T_{e}$.
2.2 Shards configuration
The shards of the cryogenic pellet after shattering are assumed to be perfect spheres in the simulations. A rejection sampling method is applied to sequentially generate the radius of each shard, presently based on Park’s fragmentation model [33]. The sampling process is terminated once the accumulated volume of the shards is equal to that of the injected pellet. The velocities of the shards are generated based on the experimental observation that the mean shard speed tends to lie between the pellet velocity parallel to the shatter plane $v_{para}$ and the injection speed of the pellet $v_{pellet}$ [21]. For this reason, we assume the fragment velocity follows a Gaussian distribution with a mean $(v_{para}+v_{pellet})/2$ and standard deviation of 20% of the mean. The directions of the shard velocities are uniformly sampled on the unit sphere within the boundaries of a circular cone whose apex is located at the center of the sphere with a $20^{\circ}$ opening angle. The starting position of all shards is the midpoint between the bend and the exit point of the shatter tube, and the average shard direction is set parallel to the line defined by these two points and going towards the plasma.
Three different shatter heads were installed on AUG during the 2021/22 experimental campaign [12, 19]. To mimic the realistic shard plume and compare the effects of different shard sizes, three different shard plumes are generated according to the above procedure, as shown in figure 1. The initial speed of the pellet is set to 443 m/s for the base scenario, with a pellet length of 10 mm and a radius of 8 mm, matching the full-sized mean pellet length around 9.6mm observed for the AUG system [19]. By varying the shattering angle from $12.5^{\circ}$ $25^{\circ}$, different shard sizes can be obtained while keeping a comparable penetration speed. The $25^{\circ}$ and $12.5^{\circ}$ cases correspond to the set-up of guide tube 1 (GT1) and GT3 in AUG, respectively. To investigate the effect of penetration speed, two additional shard configurations were created by maintaining the same shard sizes as in the 443 m/s cases with 53 and 199 shards, respectively, while manually reducing the velocities of the shards to half of their original values. All the shard configurations used in this study are listed in Tab. 1. In the simulation, each shard has its own velocity and radius, and follows an individual trajectory originating from its initial location. The velocity vector is kept constant while the radius is updated according to the ablation of the shard on its course through the plasma domain (rocket effect not taken into account).
Figure 3: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity, (d) ratio of assimilated neon to total injected neon and (e) radiation power for the 53 shard cases with different neon fraction. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
Figure 4: Temporal evolution of total radiated energy, ohmic heating energy and total thermal energy for the 53 shard cases with different neon fraction. The dotted line shows the real-time radiation fraction $frad(t)=W_{rad}(t)/[W_{th}(0)-W_{th}(t)+W_{ohm}(t)]$. The percentages in the plot inform the radiation fraction before and after 1 ms (marked by vertical lines), i.e. 0 to 1 ms and 1 ms to the end of simulation. The radiation fraction before 1 ms is equal to $frad(1~{}\mathrm{ms})$, while the radiation fraction after 1 ms is calculated by $[W_{rad}(\mathrm{end})-W_{rad}(\mathrm{1~{}ms})]/[W_{th}(\mathrm{1~{}ms})-W_{th}(\mathrm{end})+W_{ohm}(\mathrm{end})-W_{ohm}(\mathrm{1~{}ms})]$.
Figure 5: Poincaré plots of magnetic fields for the pure deuterium injection with 53 shards. The black markers indicate the real-time location of the shards.
Figure 6: (a)Perturbed poloidal magnetic flux $\delta\psi$ and (b)Poincaré plot for the same case as figure 5 at 0.57 ms. Only n=0 and n=1 toroidal harmonics of field lines are traced here to visualize the underlying island structures behind the stochastic fields.
Figure 7: Electron temperature and density at different times (a) 0.479 ms (b) 0.497 ms (c) 0.512 ms (d) 0.540 ms during the inital thermal energy drop phase for the pure deuterium case with 53 shards. The white markers indicate the real-time location of the injected shards. The white contour illustrates the q = 3 resonant surface.
Figure 8: Flux averaged temperature profile evolution for different neon fraction during the first stage cooling.
Figure 9: Magnetic energy spectrum evolution for the 53 shard cases with different neon fraction.
Figure 10: Flux averaged temperature profile evolution for the 53 shard cases with different neon fraction. The dashed lines indicate the estimated shard trajectory.
3 Numerical results
3.1 Simulation set-up
The equilibrium profiles, shown in figure 2, are based on the reconstruction of AUG H-mode discharge #40355 with CLISTE [34]. The toroidal field $B_{t}$ is 1.8 T, and the plasma current $I_{p}$ is 0.8 MA. A Spitzer-like resistivity $\eta$ is adopted in the simulation:
$\eta=\eta_{0}T_{e}^{-3/2}\cdot Z_{\mathrm{eff}}\frac{1+1.198Z_{\mathrm{eff}}+0.222Z_{\mathrm{eff}}^{2}}{1+2.966Z_{\mathrm{eff}}+0.753Z_{\mathrm{eff}}^{2}},$ (8)
where $Z_{\mathrm{eff}}$ is the effective charge,
$Z_{\mathrm{eff}}\equiv\frac{\sum_{i}n_{i}Z_{i}^{2}}{\sum_{i}n_{i}Z_{i}}.$ (9)
In order to prevent extremely large $\eta$ values, a temperature threshold $T_{min}$ = 1 eV is used to cut off the temperature dependency, i.e., when the temperature drops below this value, $\eta$ will not further increase. In the same vein, an upper limit of temperature dependency is applied at $T_{max}$ = 1.35 keV, corresponding to a $\eta$ floor value of $2~{}\times~{}10^{-8}~{}\Omega$m to simplify the simulation numerically. The perpendicular viscosity $\mu$ and parallel viscosity $\mu_{\parallel}$ are both set as $4.85\times 10^{-7}~{}\mathrm{kg~{}m^{-1}s^{-1}}$. The corresponding kinematic viscosity in the center $\nu_{0}=\mu/\rho_{0}$ is $1.6\ \mathrm{m^{2}/s}$. The perpendicular thermal diffusivity is set as $\chi_{\perp}=1\ \mathrm{m^{2}/s}$. Spitzer-Haerm parallel thermal diffusivity,
$\chi_{\|,SH}=3.6\cdot 10^{29}\frac{T_{e}[\mathrm{keV}]^{5/2}}{n_{e}\left[m^{-3}\right]}m^{2}/\mathrm{s},$ (10)
is adopted in the parallel direction. A floor value at 50 eV is applied in the simulation to help deal with low temperature regions such as temperature holes. The isotropic particle diffusivity for main ions $D$ is set to be 1.6 $\mathrm{m^{2}/s}$. The impurity diffusivity is set to be the same with the main ions, i.e. $D_{imp}=1.6\ \mathrm{m^{2}/s}$. No background impurity is included in all the simulations to keep the focus on the injected material.
The simulation utilizes a flux-aligned grid with a resolution of 70 $\times$ 110 in the radial and poloidal direction, respectively. 8 toroidal harmonics n = 0…7 are included in our simulations. Since the simulations are aimed for late TQ to early CQ, an ideal conducting wall is assumed.
The SPI is launched at the very beginning in our simulations. Once a shard enters the plasma domain, the ablation rate for a mixed neon and deuterium shard is calculated based on an improved version [35] of the neutral gas shielding (NGS) model [36, 37]. The ablation rate of each shard is obtained by
$G[g/s]=$ $\ [27.08+\mathrm{tan}(1.49X)]\left(\frac{T_{e}[\mathrm{eV}]}{2000}\right)^{5/3}$ (11)
$\ \left(\frac{r_{p}[\mathrm{cm}]}{0.2}\right)^{4/3}n_{e}[10^{14}\mathrm{cm}^{-3}]^{1/3},$
where $X=N_{D_{2}}/(N_{D_{2}}+N_{Ne})$ is the molecular mixture ratio of deuterium and $r_{p}$ is the radius of the shard. The ablated material is assumed to be a Gaussian shape source term centered at the shard location, with the following form:
$S_{n}\propto$ $\exp\left(-\frac{\left(R-R_{\mathrm{s}}\right)^{2}+\left(Z-Z_{\mathrm{s}}\right)^{2}}{\Delta r_{\mathrm{NG}}^{2}}\right)\quad$ (12)
×$\times\exp\left(-\left(\frac{\phi-\phi_{\mathrm{s}}}{\Delta\phi_{\mathrm{NG}}}\right)^{2}\right).$
In equation (12), $R_{\mathrm{s}},Z_{\mathrm{s}}$ and $\phi_{\mathrm{s}}$ are the real-time location of one specific shard. $\Delta r_{\mathrm{NG}}$ is the radius of the neutral gas cloud in the poloidal plane, which is set to 1.4×$\times r_{p}$ with a minimum value of 0.03 m, where $r_{p}$ is the radius of the shard. $\Delta\phi_{\mathrm{NG}}$ is the toroidal extent of the cloud, which is set to 0.8 rad in our simulation.
3.2 Neon content scan
To investigate the effect of the neon concentration, a scan in the neon content is performed at fixed number of fragments (LF_FV_Ne0, LF_FV_Ne0.12, LF_FV_Ne1 and LF_FV_Ne10 in Tab. 1). To isolate the effects of the different neon fractions, we assume an identical size and velocity of the shards in the scan. The total injected neon and deuterium atoms are listed in Tab. 1.
Figure 3 shows the evolution of global quantities over time after the injection. From the thermal energy evolution in figure 3 (a), it can be seen that a very rapid drop of thermal energy is triggered upon the shards arriving at the plasma domain at about 0.45 ms, regardless of the neon fraction. The sudden drop of about 50% of the thermal energy is mainly due to the convective and conductive transport caused by the magnetic field line stochastization at the edge. For the pure deuterium case (0% neon), the plasma enters a prolonged non-disruptive phase following the initial drop in thermal energy. In contrast, for neon-doped pellet cases, the thermal energy is gradually depleted after the initial drop. The pre-TQ+TQ duration (period from shards arrival to the end of TQ) increases as the neon fraction decreases. The time of shard arrival and exit are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines in the plot when the total thermal energy drops to about 100 kJ, which is in line with the the radiation peaks in figure 3 (e).
In figure 3 (c), the total assimilated neon quantity is plotted, while the assimilation ratio of impurity neon, calculated by the ratio of assimilated neon to the total injected neon content, is shown in figure 3 (d). The 10% neon pellet has the largest total assimilated neon quantity, whereas its assimilation ratio is the lowest comparing to other neon fraction cases. This is a self-regulating process that, with a lower neon fraction, the electron temperature decreases more gradually, resulting in an extended ablation period and consequently a higher assimilation ratio, which is consistent with the INDEX results in Ref. [17]. It should be noted that for the 10% and 1% neon cases, ablation stops at approximately 2 and 3 ms, respectively, when the total thermal energy drops to nearly 100 kJ, in line with the black vertical dashed lines annotating the end of TQ. However, for the 0.12% neon case, ablation saturates at around 3.5 ms as marked by the cyan vertical dashed line, while the plasma temperature remains sufficiently high for shard ablation. This occurs because, at that moment, the majority of the shards have traversed the plasma and start to exit the plasma domain. It is worth noting that even for a pellet with a low neon fraction of 0.12%, the ablation ratio remains limited to approximately 20%, indicating that a significant amount of material still remains unablated. Further increasing the neon fraction to 20% would not enhance the cooling effect of the pellet much, due to the limitation imposed by ablation [38].
Figure 3 (e) shows the evolution of radiation power versus time, where double radiation peaks are observed for the higher neon fraction cases (1% and 10%). Analysis of the radiation distribution reveals that the first radiation peak originates from the edge near the injection location, while the second peak is associated with the plasma core. The first radiation peak is due to the initial ablation of the shards, while the second peak is related to the 1/1 MHD mode during core collapse.
To understand the two-stage cooling observed in our simulations, figure 4 displays the evolution of total radiated energy, ohmic heating energy, total thermal energy and radiation fraction for the cases with 0.12%, 1% and 10% neon. It shows that, before 1 ms (indicated by the gray dashed line), the radiation fraction is very limited. In the cases with 0.12%, 1%, and 10% neon, the radiation fractions at that time are approximately 0%, 6%, and 27%, respectively. During the second stage of thermal energy drop (after 1 ms), energy loss is dominated by radiative cooling, with the radiation fraction increasing from 80% to 97% as the neon content increases. Here, the radiation fraction is calculated using the formula $\Delta W_{rad}/(\Delta W_{th}+\Delta W_{ohm})$, where $\Delta W_{rad}$ represents the change in radiated energy, $\Delta W_{th}$ is the change in thermal energy, and $\Delta W_{ohm}$ refers to the change in Ohmic heating energy. The calculation is separated into two distinct time windows of the thermal energy drop. Due to the absence of background impurity and lack of impurity radiation, the second phase may not occur at all in pure deuterium injection, like also observed in many non-disruptive shots in experiments [20].
To study the first stage of thermal energy loss in more depth, the Poincaré plots of magnetic fields for the pure deuterium injection case are plotted in figure 5. It can be seen that right after the vanguard shards reach the plasma edge (0.48 - 0.52 ms), the magnetic field becomes stochastic near the shard locations, due to the destabilization by helical cooling. Since the injection targets a H-mode plasma, the large current density gradient and pressure gradient in the pedestal are prone to MHD instabilities, triggering magnetic reconnection and leading to the stochastization. Due to the steep pressure gradient at the pedestal, consequently, considerable energy loss occurs during the early injection phase once confinement at edge is broken. Afterwards, when shards reach the q = 2 surface at around t = 0.57 ms, a large 2/1 island is induced and triggers the global reconnection event, resulting in full stochastisation of the flux surface. This large 2/1 island can be seen by producing Poincaré plots, which include only the n=0 and n=1 toroidal harmonics of the magnetic field, as shown in figure 6 (b). In figure 6 (a), the perturbed poloidal magnetic flux $\delta\psi$ at the same time is plotted, revealing a 2/1 tearing mode structure.
The evolution of electron temperature and density during first stage of thermal energy loss are shown in figure 7, and the flux averaged electron temperature profiles are plotted in figure 8. Upon the vanguard shards arriving at the q = 3 surface at 0.479 ms, the 3/1 magnetic islands are induced, as the m = 3 structure of the ablated electron density becomes clearly visible in figure 7 (a). The pedestal is subsequently degraded, and the temperature becomes flattened near the q = 3 resonant surface at around t = 0.5 ms, as shown in figure 7 (b) and in the flux averaged profile in figure 8. Following the pedestal collapse, significant heat flux propagates outward to the edge, causing a sudden ablation of the subsequent shards. This is illustrated in figure 7(c) and (d), where considerable thermal energy is expelled from the edge. Figure 8 also indicates that the majority of thermal energy loss during the first stage occurs between 0.5 and 0.6 ms. The evolution of the temperature profile is similar across different neon fractions, due to the absence of significant radiative cooling at this stage.
During the first stage of the cooling, the majority of the shards are located at the plasma edge. Despite significant degradation, the plasma temperature is still warm enough ($\sim$ 500 eV) for ablation to continue. With impurity radiation coming into effect, the second stage of cooling diverges depending on the neon fraction. The magnetic energy spectrum is compared in figure 9, for the four cases. In the pure deuterium case, only a single MHD peak is observed during the initial injection stage, corresponding to the rapid drop in thermal energy at the beginning. For neon-doped injections, in addition to the first MHD peak observed at the early stage, a later MHD peak appears at the end of the TQ phase, aligning with the radiation peaks. The second MHD peak corresponds to a 1/1 mode activity, which leads to the core collapse. In cases with a higher neon fraction, when the shards reach the core, the core collapse is triggered shortly due to stronger MHD response. With low neon concentration, the radiative cooling is less intense, leading to milder MHD activity. As a result, the plasma remains more stable for a longer period, and the core collapse occurs after the shards have exited the plasma domain.
This slow TQ process can also be seen in figure 10, where the temporal evolution of the flux averaged temperature profiles are compared for different neon fractions. For all these cases, after the initial degradation of temperature, an inside-out cooling pattern is observed. As the shards penetrate inward, they create a localized low-temperature region that moves inward along with the shards, while a temperature layer of approximately 100 eV remains at the edge. This inside-out cooling does not occur at higher shard numbers or lower penetration speed, as discussed later. For pure deuterium injection, due to lack of radiation, the thermal energy is maintained and the temperature profile gradually relaxes over time. In contrast, with neon mixed injections, the core temperature suddenly drops to below 10 eV following the 1/1 MHD activity, and then initiating the inside-out cooling.
3.3 Influence of shard size
To understand the influence of shard size on the disruption mitigation effectiveness, simulations with different shard sizes are conducted with a neon fraction of 0.12% and 10%, corresponding to the cases MF_FV_Ne0.12, SF_FV_Ne0.12, MF_FV_Ne10 and SF_FV_Ne10 from Tab. 1.
Figure 11: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different shard numbers for fixed total volume/material with 0.12% neon. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
Figure 12: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different shard numbers for fixed total volume/material with 10% neon. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
Figure 13: Flux averaged temperature profile evolution for different shard numbers with 0.12% neon. The dashed lines indicate the estimated shard trajectory.
Figure 14: Flux averaged temperature profile evolution for different shard numbers with 10% neon. The dashed lines indicate the estimated shard trajectory.
Figure 11 shows the overview plots for different shard numbers in the 0.12% neon scenario, where the total injected material is identical across cases. In figure 11 (a), a trend is observed that a larger number of smaller shards can lead to a shorter TQ duration as indicated by the thermal energy evolution. This agrees with the lower dimensional results we previously obtained between JOREK and INDEX [16]. The larger total surface area for smaller shards contributes to stronger assimilation until the majority of the shards exit the plasma at around 3.5 ms. Since the end of ablation is not determined by a full thermal collapse here, but by the leaving of the shards, the ablation for all cases saturate at almost the same time, which can be clearly seen in figure 11 (c). As the ablation duration is not constrained by the low temperature at TQ, smaller shards achieve higher assimilation compared to larger ones. Owing to the larger assimilation, the core collapse occurs earlier with smaller shards and stronger radiation power is observed at the core collapse phase, as shown in figure 11 (d).
Our numerical results deviate from the experimental findings reported in [39], where larger shards with low neon fraction exhibit better material assimilation. The key reason for this discrepancy is that our current modelling fails to accurately simulate the plasmoid drift during the pre-TQ phase and dose not take into account the rocket effect, which accelerates the shard and deflects its trajectory. Smaller shards can drive stronger plasmoid drift due to quicker ablation and larger pressure peaking, and are more susceptible to the rocket effect due to smaller mass. In experiments [39], smaller shards tend to have longer pre-TQ duration yet achieve lower assimilation. Simulations using the NGS model indicate that smaller shards ablate at a higher rate; if their total assimilation remains lower despite longer ablation durations, it is very likely due to the plasmoid drift and the rocket effect expelling the material from the plasma for low neon fraction injections.
Figure 12 shows the same data, but for the 10% neon cases with different shard numbers. It can be observed that a larger number of smaller shards leads to quicker assimilation due to the larger total surface area at the beginning, similar to cases with lower neon content. However, a reduced number of larger shards has a longer ablation duration, resulting in a relatively similar total assimilation level when ablation ceases. In contrast to the 0.12% neon cases, the ablation stops due to the temperature drop during the TQ, which occurs before the shards leave the plasma on the high field side. By the end of the TQ phase, the 199 shards case shows the highest assimilation, followed by the 1105 shards case, with the 53 shards case having the lowest assimilation. However, the trend is now less pronounced, and the assimilation levels are not significantly different, aligning with experimental observations that the size effect on assimilation is relatively small for pellets with higher neon fractions [39, 20]. This reaffirms our previous hypothesis, since the plasmoid drift is suppressed at high neon fractions. It should be noted that, in those 10% neon cases, a brief decline in assimilation is observed after the TQ. This is due to a transient increase in impurity diffusivity to 4 - 6 times its original value, i.e. $6.4-9.6\ \mathrm{m^{2}/s}$, to overcome numerical difficulties. Once this challenging period passes, the impurity diffusivity is restored to its original value. For the 199 shards case with 10% neon, the simulation could not proceed beyond the onset of the core collapse due to numerical issues.
The temperature profile evolution for the 0.12% neon cases is shown in figure 13, and for the 10% neon cases in figure 14. The 10% neon case with 199 shards is excluded from the comparison, as this simulation did not complete the TQ. In the low neon fraction cases, an inside-out cooling pattern is observed across all configurations, owing to limited impurity assimilation at the edge. However, in the high neon fraction cases, as the number of shards increases, the cooling pattern transitions from inside-out to outside-in. The plume generated by the smaller shards is more elongated, ensuring a continuous presence of shards at the edge. Additionally, smaller shards exhibit more efficient assimilation during the initial stage. Consequently, a cold front moving alongside the pellets inwards is established more effectively with smaller shards.
3.4 Effect of penetration speed
To further study the effect of penetration speed on disruption mitigation, simulations are carried out based on the cases LF_HV_Ne10 and MF_HV_Ne10 as described in Tab. 1. Generally speaking, the shard configuration is determined by two main factors: the impact speed ($v_{\perp}$), which is the pellet velocity perpendicular to the shattering plane and influences the size of the shards, and the penetration speed ($v_{\parallel}$), which is the pellet velocity parallel to the shattering plane and affects the velocity distribution of the shards after shattering. To isolate the effects of these two factors, we maintain the same shard numbers and sizes as in the cases LF_FV_Ne10 and MF_FV_Ne10 while reducing the velocities of the shards to half of their original values (Half Velocity). This approach retains the impact speed while altering the penetration speed.
Figure 15: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for the 53 and 199 shard cases with different penetration speed and 10% neon. The traces of the two slower cases are plotted with the time shifted to align the arrival time, indicated by $t_{arr}$, of the shards with that of the faster cases.
Figure 16: Flux averaged temperature profile evolution for different penetration speed with 10% neon. The dashed lines indicate the estimated shard trajectory.
Figure 17: (a) Plasma current $I_{p}$ and (b) thermal energy $W_{th}$ evolution for the case LF_HV_Ne10 in comparison with the experiment discharge #40673. In this particular experimental discharge, the pellet speed is 221 m/s, the shatter angle is 25 degrees, and the neon fraction is 10%. The traces of the simulation are plotted with the time shifted to align the arrival time of the shards with experiment, which is annotated by the cyan dotted line. The orange and blue dotted lines indicate the time of $I_{p}$ dip for the two cases, respectively.
The macroscopic quantities for different penetration speeds are compared in figure 15. The traces of the two slower cases are plotted with the time shifted to align the arrival time of the shards with that of the faster cases, for a clearer comparison. Numerical results suggest that faster penetration speeds can facilitate the material assimilation, especially during the early stage after injection, leading to shorter TQ duration. With a faster penetration speed, the shards reach the hotter regions more quickly enhancing stronger ablation, and hence increasing the impurity radiation. The initial thermal energy drop is also more steep, comparing the slow speed cases. The transition from an inside-out to an outside-in cooling pattern can be observed with decreasing penetration speed, as shown in figure 16. Due to the slower penetration speed, the shards remain longer at the edge during their penetration, allowing for more efficient impurity assimilation at the edge, which contributes to the formation of the cold front.
As the cold front forms at the edge, a more pronounced $I_{p}$ spike is observed in the simulation. Figure 17 presents the plasma current and thermal energy evolution for the case LF_HV_Ne10, compared with the experiment discharge #40673. A comparable $I_{p}$ spike with experiment is observed in the simulation. On the other hand, the $I_{p}$ dip in the simulation occurs significantly earlier than that in experiment. This earlier dip is attributed to the excessively sharp thermal energy drop following the shard arrival, as shown in the thermal energy evolution. This rapid energy loss leads to an unrealistically short pre-TQ duration, as discussed in section 3.2.
3.5 General figures of merit
The aforementioned simulations provide valuable insights into the plasma response to various shard parameters. In this subsection, general trends and summary plots are presented, highlighting the dependence on these parameters and how the results align with experimental observations.
Figure 18: Radiation fraction for all neon doped cases calculated by $f_{rad}=W_{rad}/(W_{th}(t=0)+W_{ohm})$ at the time point when the plasma current has dropped to $I_{p}=0.8I_{p}(t=0)$. The pure deuterium case is excluded, as neutral line radiation and background impurities are not yet considered in our current modelling.
Figure 19: Radiation fraction for all neon doped cases calculated by $f_{rad}=W_{rad}/(W_{th}(t=0)+W_{ohm})$ at the end of simulation, comparing with experimental results [20]. The pure deuterium case is excluded, as neutral line radiation and background impurities are not yet considered in our current modelling.
Figure 20: The total radiation power vs. toroidal peaking factor. The TPF is calculated by $P_{rad}^{max}/\Sigma(P_{rad,i}+P_{rad,i+1})/2*\Delta_{\phi}(i,i+1)/360$, where $P_{rad}^{max}$ is the maximum radiation among the 5 sectors installed with foil bolometer, $(P_{rad,i}+P_{rad,i+1})/2$ is the average radiation between the neighboring sectors and $\Delta_{\phi}(i,i+1)$ is the toroidal angle between the two neighboring sectors. The contour plot represents the constant values of the product TPF * $P_{rad}$.
Figure 21: Evolution of radiation power from different sectors for (a) the case LF_FV_Ne10 and its (b) counterpart in experiments. S16 is the injection sector and s1 and s15 are the neighboring sectors. S5 and S9 are sectors perpendicular and opposite to the injection sector, respectively.
Figure 22: Pre-TQ+TQ duration, period from shard arrival to the end of TQ, for all neon doped cases. For the pure deuterium case, since a complete TQ does not occur, this duration is not available.
Figure 23: Pre-TQ duration, period from first light $t_{FL}$ to the $I_{p}$ dip (proxy for the TQ onset), in AUG experiments.
Figure 24: Total assimilated neon atoms for all neon doped cases with different neon fractions, fragment sizes and speeds.
Figure 25: Total assimilated neon and deuterium atoms for all the cases with different neon fractions, fragment sizes and speeds.
Figure 18 illustrates the radiation fraction similar to the experimental definition following Ref. [20] $f_{rad}=W_{rad}/(W_{th}(t=0)+W_{ohm})$ at the time point when the plasma current $I_{p}$ has dropped to 80% of its original value for a consistent comparison, since not all the cases reach at the same CQ level by the end of simulation. Here, $W_{rad}$ is the total radiated energy, $W_{th}(t=0)$ is the thermal energy at the beginning, and $W_{ohm}$ is the total ohmic heating energy in the considered time interval. It should be mentioned that, in our current modeling, as described in section 2, the deuterium neutral fluid and neutral line radiation are not considered. Additionally, no background impurities are included in the 3D simulations. Consequently, for the pure deuterium case, the radiation fraction is 0%. As for neon-doped pellets, the radiated energy accounts for approximately 40% of the lost energy for the 0.12% neon case, and 45% for the 1% neon case. Furthermore, the size of the shards does not have any significant impact on the radiation fraction in the 0.12% neon cases, owing to the low radiative cooling effect during the first stage of thermal energy loss. For 10% neon cases, the radiation fraction ranges from 56% - 74%, depending on different injection scenarios. Faster and smaller shards exhibit better performance in terms of radiation fraction, as they generate higher radiation during the first stage of thermal energy loss.
To compare with experiments, the radiation fraction, calculated using the above definition at the end of each simulation, is plotted alongside experimental data in figure 19. The experimental data is derived until the end of the CQ [20]. The overall radiation fraction in the simulations is relatively lower than that observed in experiments. This discrepancy is primarily due to excessive thermal energy loss during the first stage of the cooling process, which is characterized by a low radiation level. The key to improve the radiation fraction lies in increasing radiative cooling during the first stage. Notably, the case with 10% neon and 1105 shards shows a comparable radiation fraction to experimental results, owing to more effective radiative coverage in the initial cooling phase.
The total radiation power versus the toroidal peaking factor (TPF) is plotted in figure 20, where the TPF is obtained using the same formula in experiments [40]. The key message here is that there are two distinct radiation peaks, which are much more prominent in the large neon fraction cases. The first peak occurs at the initial injection and the second at the end of the TQ. The first peak, characterized by a relatively large TPF, is mainly caused by the outward heat flux that result in sharp ablation of the shards during the first stage of thermal energy drop, which is overestimated in simulations and not observed in experiments. The smaller shard case exhibits a stronger TPF and higher radiation at this stage. The TPF at the TQ is generally very small for all cases, as the ablated material has spread spatially at the latter phase, which is also observed experimentally. It is crucial to prevent high toroidal peaking during the initial injection phase in experiments, as simulations suggest that this phase is more likely to produce a high TPF, despite the fact that our current simulations tend to overestimate the TPF at this stage. One possible explanation for the more peaked early-phase radiation peaking observed in our simulations is opacity. During the early phase, the local density around the shards could become very high so that the plasma is not transparent for line radiation, reducing the local radiation power density around the shards in reality [41, 42].
In addition, the distribution of radiation across different toroidal sectors in the experiment is more uniform (except S9) compared to what observed in simulation, as illustrated in figure 21. In the simulation, the difference in radiation between sectors near the injection plane and those far from it is more significant, comparing with experiments [40]. This may be due to the fact that the equilibrium toroidal flow is not considered in the simulation.
Figure 22 presents the pre-TQ+TQ duration for all neon-doped cases. In the case of pure deuterium injection, no complete TQ is observed due to the lack of radiative cooling, resulting in an effectively infinite TQ duration. This behavior aligns with non-disruptive shots or those with extended pre-TQ durations seen in experiments following pure deuterium injection [20]. For neon-doped injections, even a slight neon fraction of 0.12% reduces the pre-TQ+TQ time to 3 – 5 ms. As the neon fraction increases to 10%, this duration further decreases to approximately 1 ms. Additionally, for cases with low neon fractions, the pre-TQ+TQ time shows a noticeable dependence on shard size, which can be attributed to the significant differences in material assimilation in our simulations.
Figure 23 illustrates the correlation between pre-TQ duration and neon fraction observed in AUG experiments. In the pure deuterium cases, the pre-TQ durations vary significantly, ranging from a few milliseconds to tens of milliseconds. However, once the neon fraction exceeds 1%, the pre-TQ durations tend to converge, with most cases falling between 3 to 5 ms. This trend also suggests that JOREK simulations presented in this article underestimate the pre-TQ duration, especially with a low neon fraction, because of the exaggerated initial thermal energy drop.
In terms of the assimilation, figure 24 exhibits the assimilated neon atoms for all the neon doped cases. As expected, neon assimilation in absolute numbers generally increases with higher neon fractions. For the 0.12% neon cases, there is a clear dependence on the shard size, with smaller shards showing higher assimilation, which deviates from the experimental results. This is because our current simulation fails in accurately simulating the plasmoid drift and does not take into account the rocket effect, as we discussed in section 3.3. On the other hand, for cases with higher neon fractions, such as the 10% neon case, the dependence on shard size is less pronounced due to the extended ablation duration for larger shard case. This part of results is consistent with experimental observation, presumably due to the suppression of plasmoid drift at high neon fraction in experiments. With respect to penetration speed, faster shards result in higher assimilation during the early stage after injection, as shown in figure 15; however, the difference is very small in the end.
Regarding the total assimilation, figure 25 displays the total assimilated atoms (neon + deuterium) for all cases. The trend reverses when considering the total assimilated atoms, with total assimilation increasing as the neon fraction decreases. This is due to the lower assimilation ratio observed in high neon fraction cases, primarily because of the much shorter TQ duration. As these pellets contain comparable amounts of deuterium, and deuterium atoms far outnumber neon atoms within the pellet, a higher ablation ratio leads to enhanced deuterium assimilation. The dependence on shard size and speed follows a similar pattern to that of neon assimilation.
4 Summary and discussion
In this study, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using realistic parameters with the 3D non-linear JOREK code, aimed at validating optimal shard configurations for the ITER DMS. The key findings can be summarized as follows:
Two distinct cooling phases are identified in the simulations. The first is dominated by convective and conductive transport caused by magnetic field stochastization at the plasma edge, virtually independent of injection parameters. The second phase, more prominent in neon-doped cases, is driven by radiative cooling, which is critical for the TQ duration. As for pure deuterium injection without background impurity, the second phase of cooling may not occur at all due to limited radiation, corresponding to non-disruptive shots observed in experiments. Results show that high neon fraction (10%) leads to faster TQ through radiative processes, while lower fraction (0.12%) result in more gradual cooling. High neon fraction injections (10%) cause stronger MHD response during the second stage cooling, leading to faster core collapse and more intense cooling. Low neon fractions slow down this process, leading to longer TQ durations and a more gradual loss of thermal energy.
The total assimilation becomes less efficient with increasing neon content. Higher neon fraction can lead to shorter ablation duration, resulting in a lower ablation ratio. For the low neon fraction (0.12%) cases, the assimilation exhibits a strong dependence on shard size. Smaller shards increase the surface area available for ablation, leading to higher assimilation. This trend diverges from experimental observations, where larger shards typically lead to higher material assimilation. This discrepancy arises from inaccuracies in simulating the pre-TQ phase. A larger number of smaller shards can induce stronger plasmoid drift due to quicker ablation and larger pressure peaking, which is not accurately captured in the current simulation. Additionally, the current model does not incorporate the rocket effect, which might be particularly important for small shards, as their smaller mass makes them more susceptible to this effect. For the cases with high neon fraction (10%), the dependence of assimilation on shard size is less pronounced. The smaller shards lead to shorter ablation duration, even though the ablation rate is higher. This part of results aligns with experimental observations, likely due to the suppression of plasmoid drift at high neon fraction injections in experiments.
In cases involving large shards (53 shards), an inside-out cooling pattern emerges due to less efficient edge cooling. The results suggest that smaller and slower shards promote edge cooling, which aids in forming a cold front and shifting the cooling pattern from inside-out to outside-in. The plume generated by smaller shards is more elongated, ensuring a continuous presence of shards at the edge. Additionally, smaller shards demonstrate more efficient assimilation during the initial stages, while slower shards remain longer at the edge during penetration, further enhancing edge cooling. As the cold front forms at the edge, an $I_{p}$ spike comparable to the experimental one is observed in the simulation.
The overall radiation fractions observed in the simulations are slightly lower than those seen in experiments, likely due to overestimated thermal energy loss in the first stage of cooling. This discrepancy is primarily due to excessive thermal energy loss during the first stage of cooling, which is not sufficiently radiated. As for the radiation asymmetry, two radiation peaks are observed in large neon fraction cases, one at initial injection and another at the end of the TQ. The first peak, caused by outward heat flux during the initial thermal energy drop, is overestimated in our simulations. The smaller shard case shows stronger peaking factors and higher radiation during this stage. The radiation peak at TQ shows a small peaking factor due to a more relaxed plasma state. The radiation distribution in experiments is more uniform across toroidal sectors than in simulations, likely because our current simulations do not include the equilibrium toroidal flow and the opacity effect. The current work aims to validate the SPI model in JOREK and compare the general trends with experimental results. More detailed comparisons with experiments will be carried out in the future work.
Acknowledgements
We sincerely thank Sang-Jun Lee, Mengdi Kong, Daniele Bonfiglio, Andres Cathey, Nina Schwarz and Haowei Zhang for fruitful discussions. We are also very grateful for the CPU time generously provided by the HAWK supercomputer operated by the High Performance Computing Center (HLRS) in Stuttgart, Germany. Part of this work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Within EUROfusion, in particular the Theory and Simulation Verification and Validation (TSVV) projects on MHD Transients (PI M. Hoelzl) and the European HPC infrastructure (Marconi-Fusion) provided essential support. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. Part of this work has been carried out in collaboration with the ITER Organization. The views and opinions expressed herein do not necessarily reflect those of the ITER Organization. This work receives funding from the ITER Organization under contract IO/20/IA/43-2200. The ASDEX-Upgrade SPI project has been implemented as part of the ITER DMS Task Force programme. The SPI system and related diagnostics have received funding from the ITER Organization under contracts IO/20/CT/43-2084, IO/20/CT/43-2115, IO/20/CT/43-2116.